OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions: A. (6x-5y=11; 10y=9x-22) B. (8x-4=6y; 12y-8=4x) C. (4x-7y=4; 14y=8x-8) D. (5y=-6x-4; 12x+15y=8)

OpenStudy (anonymous):

they represent st.lines.......they infinte solutions if they overlap.......it means they shud represent the same line...........now can u try?

OpenStudy (anonymous):

i have no idea what you mean

OpenStudy (anonymous):

ok check c 14y=8x-8 8x-14y=8 2(4x-7y)=2(4)

OpenStudy (anonymous):

@lalaly

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

choose the one that when you change both equations into y=mx+b form, if they're exactly the same, that' the one you want... solve for y in each choice... which one is it?

OpenStudy (anonymous):

@dpaInc : C ?

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

yw...:)